# references.bib

### v1

@@ 1 Introduction
    the Square Root Law is in a first approximation independent of the time $T$ needed to execute the metaorder, and only depends on the volatility of the asset and the fraction of the total traded volume captured by $Q$ [bouchaud2018trades, webster2023handbook];
@@
  • Section 3 presents several evidences of the method effectiveness in replicating and validating the well-documented empirical facts about metaorders.
@@ 2 The Algorithm
In this section, we present our algorithm designed to generate random metaorders from publicly available data, which lead to impact properties indistinguishable from actual proprietary data. We define a metaorder as a sequence of trades of the same sign initiated by a given trader within the same trading session. In order to generate synthetic metaorders, we propose the following algorithm, which requires only public trade data for a specific stock. Although using aggregated order book data across multiple venues yields similar results, we recommend using data from a single exchange (ex : Euronext, Nasdaq, CME etc…).
@@ 3.1 Peak impact: the Square Root Law
We first tested our algorithm on a very liquid asset: the EuroStoxx futures contract from September 2016 to August 2018. We used public trade data, selecting only the Eurex exchange. We obtain a remarkably clean square root law over four decades, see Figure 1, with a noisy region for very small $Q/V_{D}\leq 5\times 10^{-6}$, which might possibly be considered as a linear, as in [Buccicrossover]. Note that we do not only recover the square-root dependence on $Q$ but also the correct prefactor in Eq. (2), with $Y=0.5$.
Figure 1: Retrieving the Square Root Law for futures on the Eurostoxx, trade data from 2016 to 2018. We used a mapping function with 20 traders and a power law distribution of their trading frequencies with an exponent of $\alpha=2$, resulting in approximately 3 million metaorders.
We also tested these results on single stocks by selecting a basket of seven stocks traded on the Paris Stock Exchange, between January 2021 and December 2023. We used our algorithm to generate approximately 3 millions of random metaorders per stock, imposing 20 traders and a homogeneous trading frequency distribution. Once again, we obtained very precise results, with almost no variation in the prefactor ($Y\approx 0.5$), see Figure 2. This could be explained by the fact that selected stocks are among the most liquid ones traded on the PSE, and thus may be quite similar. We also tested for a specific stock (BNP Paribas) the dependence of the SQL on the input parameters, i.e. the number of traders and the distribution of their trading frequencies. Again, we found no significant variations in the impact function. However, we do acknowledge that for some assets, particularly illiquid ones, one may have to fine-tune those parameters to recover the SQL. It is necessary to be statistically close enough to the real mapping function.
@@ 3.3 Concave profile during metaorder execution
Such a concave profile can be explained within the latent liquidity framework, which predicts that the latent limit order book is locally linear [donier2015fullyconsistentminimalmodel, bouchaud2018trades]. The fact that this profile also holds for synthetic metaorders is another indication that, on average, the latent order book is indeed present and provides liquidity for all metaorders. This could be related to the action of market makers, as proposed in [maitrier2025double]. That said, the concavity is in fact crucial for market efficiency : it may be a key element to ensure the diffusivity of price as argued recently in [sato2025exactlysolvablemodelsquareroot]. However, we believe that the framework developed in that paper is insufficient to accurately describe reality, as it lacks a crucial component: metaorder decay, to which we turn next. A unified theory of price impact that incorporate all known ingredients (autocorrelation of the sign of the trades, square-root impact, impact decay) is still under construction, a topic on which we hope to report soon [ustocome].
@@ 3.4 Metaorder decay post execution
Impact decay has been subject to controversy, even when it is of crucial importance for optimal execution schedules. Indeed, assuming permanent impact or accounting for impact decay leads to radically different trading policies. What makes the empirical study of this problem particularly difficult is the fact that price variance increases linearly with the time elapsed since the end of execution, leading to large error in the determination of impact decay. We know that metaorder impact during execution is generally small relative to the volatility (see Eq. (2) for small $Q/V_{D}$), this predicament is especially strong for metaorder decay: the signal-to-noise ratio significantly worsens when analyzing extended timescales after execution, highlighting the need for large metaorder datasets.
@@
Figure 5: Price relaxation post metaorder execution, well fitted by the impact predicted by the propagator model, with the value $\beta\approx 0.2$. We used synthetic metaorders generated on BNP Paribas share price between 2021 and 2023, using a mapping function of parameters of 4 traders and power law distribution of exponent $\alpha=2$. We kept only metaorders with more than 5 childs.
@@ Acknowledgments
We would like to thank Natascha Hey, Kiyoshi Kanazawa, Fabrizio Lillo, James Ridgway, Yuki Sato and Bence Toth for helpful discussions and suggestions. This research was conducted within the Econophysics & Complex Systems Research Chair, under the aegis of the Fondation du Risque, the Fondation de l’École Polytechnique and Capital Fund Management.

### v2

@@ 1 Introduction
    the SQL is in a first approximation independent of the time $T$ needed to execute the metaorder, and only depends on the volatility of the asset and the fraction of the total traded volume captured by $Q$ [bouchaud2018trades, webster2023handbook];
@@
  • Section 3 presents several evidences of the method’s effectiveness in replicating and validating the well-documented empirical facts about metaorders.
@@ 2 The Algorithm
In this section, we present our algorithm designed to generate random metaorders from publicly available data, which lead to impact properties indistinguishable from actual proprietary data. We define a metaorder as a sequence of trades of the same sign initiated by a given trader within the same trading session. In order to generate synthetic metaorders, we propose the following algorithm, which requires only public trade data for any asset class (stock, futurs, options etc…). Although using aggregated order book data across multiple venues yields similar results, we recommend using data from a single exchange (ex : Euronext, Nasdaq, CME etc…).
@@
Note that a key aspect of this mapping function is that it corresponds to sampling without replacement. We find this feature essential for recovering the SQL.
@@ 3.1 Peak impact: the Square Root Law
We first tested our algorithm on a very liquid asset: the EuroStoxx futures contract from September 2016 to August 2018. We used public trade data, selecting only the Eurex exchange. We obtain a remarkably clean square root law over four decades, see Figure 1, with a noisy region for very small $Q/V_{D}\leq 5\times 10^{-6}$, which might possibly be considered as a linear, as in [Buccicrossover]. Note that we do not only recover the square-root dependence on $Q$ but also the correct a realistic prefactor in Eq. (2), with $Y=0.5$, as found in [toth2011anomalous, sato2024does].
Figure 1: Retrieving the Square Root Law for futures on the Eurostoxx, trade data from 2016 to 2018. We used a mapping function with 4 trades (resp. 40 trades) for the blue curve (resp. orange), and homogeneous distribution of their trading frequencies, resulting in approximately 3 million metaorders in both cases. We see that, by fine-tuning the number of traders, one may recover the correct prefactor $Y\approx 0.5$
We also tested these results on single stocks by selecting a basket of seven stocks traded on the Paris Stock Exchange, between January 2021 and December 2023. We used our algorithm to generate approximately 3 millions of random metaorders per stock, imposing 20 traders and a homogeneous trading frequency distribution. Once again, we obtained very precise results, with almost no variation in the prefactor ($Y\approx 0.5$), see Figure 2. This could be explained by the fact that selected stocks are among the most liquid ones traded on the PSE, and thus may be quite similar. We also tested for a specific stock (BNP Paribas) the dependence of the SQL on the input parameters, i.e. the number of traders and the distribution of their trading frequencies. Again, we found no significant variations in the impact function. However, we do acknowledge that for some assets, particularly illiquid ones, one may have to fine-tune those parameters to recover the SQL. It is necessary to be statistically close enough to the real mapping function. The same remark also applies to the Y-ratio. While we mostly show impact functions with a realistic prefactor, i.e., close to 0.5, we have also encountered stocks where the prefactor was slightly higher of lower, but typically of order one, see Figure 1.
@@ 3.2 Role of metaorder duration
Figure 3 also reveals that, on average, synthetic metaorders constructed using our method are shorter than those typically found in proprietary datasets. For instance, it is not uncommon for firms like CFM to execute metaorders over an entire trading day. However, the durations of our synthetic metaorders are consistent with the average metaorder duration observed in the TSE dataset, which is typically also around 2-5 minutes, see [maitrier2025double]. In any case, one can adjust metaorder duration by tuning the mapping function’s parameters.
@@ 3.3 Concave profile during metaorder execution
Such a concave profile can be explained within the latent liquidity framework, which predicts that the latent limit order book is locally linear (LLOB) [donier2015fullyconsistentminimalmodel, bouchaud2018trades]. The fact that this profile also holds for synthetic metaorders is another indication that, on average, the latent order book is indeed present and provides liquidity for all metaorders. This could be related to the action of market makers, as proposed in [maitrier2025double]. That said, the concavity is in fact crucial for market efficiency : it may be a key element to ensure the diffusivity of prices as argued recently in [sato2025exactlysolvablemodelsquareroot]. However, we believe that the framework developed in that paper is insufficient to accurately describe reality, as it lacks a crucial component: metaorder decay, to which we turn next. A unified theory of price impact that incorporate all known ingredients (autocorrelation of the sign of the trades, square-root impact, impact decay) is still under construction, a topic on which we hope to report soon [ustocome].
@@ 3.4 Metaorder decay post execution
Impact decay has been subject to controversy, even when it is of crucial importance for optimal execution schedules. Indeed, assuming permanent impact or accounting for impact decay leads to radically different trading policies. What makes the empirical study of this problem particularly difficult is the fact that price variance increases linearly with the time elapsed since the end of execution, leading to large errors in the determination of impact decay. We know that metaorder impact during execution is generally small relative to the volatility (see Eq. (2) for small $Q/V_{D}$), this predicament is especially strong for metaorder decay: the signal-to-noise ratio significantly worsens when analyzing extended timescales after execution, highlighting the need for large metaorder datasets.
@@
Figure 5: Price relaxation post metaorder execution, well fitted by the impact predicted by the propagator model, with the value $\beta\approx 0.2$. We used synthetic metaorders generated on BNP Paribas share price between 2021 and 2023, using a mapping function of parameters of 4 traders and power law distribution of exponent $\alpha=2$. We kept only metaorders with more than 5 child orders.
@@ Acknowledgments
We would like to thank Doyne Farmer, Natascha Hey, Kiyoshi Kanazawa, Fabrizio Lillo, James Ridgway, Yuki Sato and Bence Toth for helpful discussions and suggestions. This research was conducted within the Econophysics & Complex Systems Research Chair, under the aegis of the Fondation du Risque, the Fondation de l’École Polytechnique and Capital Fund Management.
@@
Synthetic metaorders constructed on synthetic prices
A natural question arising from this empirical study is whether public trade data are actually needed to construct synthetic metaorders that verify the SQL. In other words, is there a specific element or piece of information in real trade data that distinguishes actual prices from random prices?
A possible experiment is to generate a basic synthetic price. This can be achieved by first creating a sequence of random volumes and random signs. Each order is then assigned a signed instantaneous impact, which can be a linear or concave function of the volume. Then, returns read :
$r_{t}=\varepsilon_{t}q^{\chi},\quad\chi\in\{0,\frac{1}{2},1\},\quad q\sim\mathcal{U}(q_{min},q_{max})$ (5)
In the following simulation we set $q_{min}=1$ and $q_{max}=100$, but the same results hold for $q=q_{\min}=q_{\max}=1$, just generating synthetic metaorders with a lower average size and variance. Based on [maitrier2025double], we fixed $\chi=0.5$.
This process produces synthetic trade data that can serve as a foundation for constructing random metaorders. However, metaorders constructed with this algorithm won’t exhibit SQL, but a linear impact function, see Figure 8.
Since trade signs are known to be autocorrelated, it is tempting to leverage this well-established stylized fact to recover the square root law. We generate a series of trade signs with an autocorrelation $C(l)=\langle\varepsilon_{t}\varepsilon_{t+l}\rangle=l^{-\gamma}$ and we naturally impose a propagator-type impact $G(t)\approx t^{-\beta},\quad\beta=\frac{1-\gamma}{2}$, see [bouchaud2018trades], to recover price diffusivity. But again, by doing so, one will still obtain a linear impact for synthetic metaorders, see Figure 8.
Figure 8: Retrieving a linear impact when constructing synthetic metaorders on a synthetic price. The basic price is built using uniformly distributed volumes ($q_{min}=1$ and $q_{max}=100$ and a series of uncorrelated random signs. The instantaneous impact reads $r_{t}=\varepsilon_{t}q^{\chi}$ with $\chi=0.5$. Synthetic metaorders were constructed with 4 traders and a power law distribution of exponent $\chi=2$.The propagator price refers to a synthetic price constructed with autocorrelated signs, $C(l)=\langle\varepsilon_{t}\varepsilon_{t+l}\rangle=l^{-\gamma}$ and a transient impact that decays according to the propagator model : $G(l)=t^{-\beta}$